Weaver Company’s predetermined overhead rate is $20.00 per direct labour-hour, and its direct labour wage rate is $15.00 per hour. The following information pertains to Job A-200:

    
  Direct materials $240
  Direct labour $180
Required:
1.

What is the total manufacturing cost assigned to Job A-200?

     

2.

If Job A-200 consists of 55 units, what is the average cost assigned to each unit included in the job?
Answer #1

Total labor hours=($180/15)=12 hours

Hence total overhead would be=($20/direct labor hour*12 hours)=$240

a.total manufacturing cost =Direct materials+Direct labor+Total overhead

=(240+180+240)=$660

b.Average cost assigned=(660/55)=12 per unit.